OpenStudy (anonymous):

What is the common difference in the sequence 8, 16, 24, 32, . . .? A. 24 B. 16 C. –8 D. 8

OpenStudy (anonymous):

d

jigglypuff314 (jigglypuff314):

16-8 = 24-16= 32-24=

OpenStudy (anonymous):

16 - 8 = 8

OpenStudy (anonymous):

D 8. just subtract the last entry from the one before it, and then repeat it like this 32 - 24 = 8 24 - 16 = 8 16 - 8 = 8

OpenStudy (anonymous):

The floor of a rectangular room will be tiled with square tiles that measure 3 inches on each side. The room is 109 inches long and 103 inches wide. How many tiles, including parts of tiles, are needed to complete the job? A. 1,248 tiles B. 1,247 tiles C. 1,224 tiles D. 1,295 tiles

OpenStudy (anonymous):

Each tile has area 3 x 3 = 9 square inches. The entire room has area, 109 x 103 = 11,227 square inches So number of tiles is 11,227/9 = approx 1,247

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

(1/2)(8^2) - (1/2)(8) = 32 - 4 = 28

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

welcome